# Inverse Problem Sampling in Latent Space Using Sequential Monte Carlo

### v1

@@ 4.2.2 Posterior Sampling
Table 1: Quantitative results on $1024$ examples of size $256\times 256$ from ImageNet test set. All methods were evaluated under the same experimental setup using LDM.
Inpainting (Box) Gaussian Deblur Super Resolution (8×8$\times$8 ×) FID ($\downarrow$) NIQE ($\downarrow$) LPIPS ($\downarrow$) FID ($\downarrow$) NIQE ($\downarrow$) LPIPS ($\downarrow$) FID ($\downarrow$) NIQE ($\downarrow$) LPIPS ($\downarrow$) Latent DPS 65.45 7.918 0.407 52.48 6.855 0.383 61.02 6.514 0.439 Latent TDS 65.03 7.872 0.406 50.82 6.695 0.379 58.73 7.157 0.454 Resample 90.32 8.464 0.318 46.45 7.411 0.353 87.65 8.290 0.491 PSLD 79.90 9.268 0.410 79.31 7.972 0.474 78.56 7.000 0.467 LD-SMC (Ours) 51.81 5.103 0.355 52.17 6.789 0.382 59.27 6.423 0.437
4.2.3 Proposal Distribution
@@
The SISR algorithm requires access to $\pi(\tilde{{\mathbf{z}}}_{t-1}|\hat{{\mathbf{z}}}_{t}^{(i)},{\mathbf{y}}_{0:T-1})$, a proposal distribution. The optimal choice in the sense of minimizing the variance of the proposal weights is $\pi(\tilde{{\mathbf{z}}}_{t-1}|\hat{{\mathbf{z}}}_{t:T}^{(i)},{\mathbf{y}}_{0:T-1})=p_{\theta}(\tilde{{\mathbf{z}}}_{t-1}|\hat{{\mathbf{z}}}_{t}^{(i)},{\mathbf{y}}_{0:t})$ (Doucet et al., 2000; Särkkä, 2013). However, it cannot be obtained in closed form. Hence, we design an alternative proposal distribution, which we will now describe.
For clarity, we drop here the index notation of the particles. We define the proposal distribution to be a Gaussian $\pi(\tilde{{\mathbf{z}}}_{t-1}|\hat{{\mathbf{z}}}_{t:T},{\mathbf{y}}_{0:T-1})=\mathcal{N}({\mathbf{m}}_{t},{\mathbf{S}}_{t})$ with parameters:
${\mathbf{S}}_{t}$ $=\sigma_{t}^{2}{\mathbf{I}}$ (4)
${\mathbf{m}}_{t}$ $=\bm{\mathbf{\mu}}_{\theta}(\hat{{\mathbf{z}}}_{t},t)-(\gamma_{t}\nabla_{\hat{{\mathbf{z}}}_{t}}||{\mathbf{y}}_{0}-{\mathcal{A}}({\mathcal{D}}(\bar{{\mathbf{z}}}_{0}(\hat{{\mathbf{z}}}_{t})))||_{2}^{2}$
$\qquad\qquad~{}+\lambda_{t}\nabla_{\bm{\mathbf{\mu}}_{\theta}(\hat{{\mathbf{z}}}_{t},t)}||{\mathbf{y}}_{t}-{\mathcal{A}}({\mathcal{D}}(\bm{\mathbf{\mu}}_{\theta}(\hat{{\mathbf{z}}}_{t},t)))||_{2}^{2}).$
@@
Here we set the variance to be the variance of the prior diffusion model; however, other choices are applicable as well. The idea behind our proposal mean is to correct the prior mean estimation by shifting it towards latents that agree more strongly with both ${\mathbf{y}}_{t}$ and ${\mathbf{y}}_{0}$. The second term can be seen as making one gradient update step starting from the current prior mean location.
@@
The $\gamma_{t}$ and $\lambda_{t}$ parameters control the effect of the correction terms to the prior mean. In practice, it is challenging to control the trade-off between the two correction terms. Hence, during the sampling process we start only with the first term and then at some predefined point, $t=s$, we switch to the second term. The intuition here is that during the initial sampling steps the quality of the labels ${\mathbf{y}}_{t}$ may not be good. Therefore, we rely on the first term through the posterior mean estimator to capture the general semantics of the image. However, in later sampling stages the quality of the labels increases (see Figure 5 in the Appendix) and the latter correction term can add fine details to the image. This intuition also relates to the three-stage phenomenon in the diffusion sampling process witnessed in the literature (Yu et al., 2023). Setting $\lambda_{t}=0$ for all time steps reduces the LD-SMC proposal update to that of TDS (Wu et al., 2024). We show in Appendix B an instantiation of $\gamma_{t}$ and $\lambda_{t}$ used in this study.
@@ 5 Experiments
Table 2: Quantitative results on $1024$ examples of size $256\times 256$ from FFHQ test set. All methods were evaluated under the same experimental setup using LDM.
Inpainting (Box) Gaussian Deblur Super Resolution (8×8$\times$8 ×) FID ($\downarrow$) NIQE ($\downarrow$) LPIPS ($\downarrow$) FID ($\downarrow$) NIQE ($\downarrow$) LPIPS ($\downarrow$) FID ($\downarrow$) NIQE ($\downarrow$) LPIPS ($\downarrow$) Latent DPS 39.81 7.592 0.236 31.81 6.813 0.285 29.64 6.412 0.282 Latent TDS 39.57 7.602 0.236 33.19 6.879 0.288 30.45 6.411 0.284 Resample 86.79 7.142 0.230 39.80 7.441 0.275 59.23 7.307 0.356 PSLD 47.51 7.480 0.312 36.31 6.802 0.341 40.33 6.803 0.347 LD-SMC (Ours) 37.14 7.520 0.224 32.18 6.566 0.280 30.37 6.456 0.284
@@ 5.1 Experimental Setting
We evaluated LD-SMC on ImageNet (Russakovsky et al., 2015) and FFHQ (Karras et al., 2019); both are common in the literature of inverse problems (e.g., (Chung et al., 2023b; Dou & Song, 2024)). In ImageNet samples were conditioned on the class label. The guidance scale was fixed to $1.0$ in all our experiments. Results can be improved by adjusting it (Rombach et al., 2022). Images were resized to $3\times 256\times 256$ and normalized to the range $[0,1]$. We used the latent diffusion model (LDM) VQ-4 (Rombach et al., 2022) for the prior model with the DDIM diffusion sampler (Song et al., 2021a), according to the data split in (Esser et al., 2021). We sampled $1024$ random images from the validation set of each dataset which were used to evaluate all methods. We followed the protocol of (Song et al., 2024) and added Gaussian noise with zero mean and standard deviation $\tau=0.01$ to the corrupted images. Full experimental details are provided in Appendix A.
@@
Compared methods. We compared LD-SMC with several recent SoTA baseline methods; the first two methods were designed for general inverse problems, but were evaluated using pixel space diffusion models, and the latter two were designed specifically for inverse problems in the latent space. Nevertheless, all methods were evaluated under a similar experimental setup using latent diffusion to ensure fairness in the comparisons. (1) Diffusion Posterior Sampling (DPS) (Chung et al., 2023b), which introduces correction to the sampling process of the diffusion through the posterior mean estimator; (2) Twisted Diffusion Sampling (TDS) (Wu et al., 2024), which uses the twisting technique for approximate sequential Monte Carlo sampling; (3) Resample (Song et al., 2024), which applies an optimization procedure during the sampling process to match the approximate posterior mean to the label, and then performs resampling; (4) Posterior Sampling with Latent Diffusion (PSLD) (Rout et al., 2023), which introduces a correction term to the DPS step to “glue” ${\mathbf{z}}_{0}$. For our method and TDS we used $N=5$ particles.
@@
Evaluation metrics. We report the following metrics in the main text, FID (Heusel et al., 2017), NIQE (Mittal et al., 2012), and LPIPS (Zhang et al., 2018). Full results with the PSNR and SSIM metrics (Wang et al., 2004) are deferred to Appendix D. The first two are considered perceptual metrics, lower values in them indicate higher perceptual quality. The other metrics are considered as distortion metrics, which quantify some discrepancy between the generated images and the ground-truth values. Since perceptual metrics and distortion metrics can be in conflict with each other (Blau & Michaeli, 2018), we put more emphasis on perceptual quality. Hence, for all methods, we performed grid search over hyper-parameters and chose the best configuration according to the FID.
@@ 5.2 Experimental Results
Quantitative results are shown in Tables 1 and 2. From the tables, LD-SMC is usually the best or second best among all the comparisons. Specifically, on inpainting where extrapolation is needed inside the box and details should be preserved outside the box, LD-SMC can greatly improve over baseline methods, improving the FID score by up to $\sim 13$ points. This property is also manifested in Figure 2 and Figure 6 in the appendix. LD-SMC manages to produce plausible reconstructions while maintaining fine details. This is in contrast to baseline methods, which mainly rely on the posterior mean approximation for reconstruction. The differences are especially highlighted on the ImageNet dataset which has more diversity in it. Also, as is clear from the figures, Resample images suffer from significant artifacts. We speculate that it partly stems from the complete optimization process performed in every few sampling steps according to this method. We observe that this method is well suited for some tasks, such as Gaussian deblurring, but does not perform well on others such as box inpainting.
In Figure 4 we show qualitative results for Gaussian deblurring of images by LD-SMC and baseline methods. From the figure, all methods are able to generate plausible reconstructions on this task. Additional quantitative and qualitative results can be found in Appendix D and E .

### v2

@@ 4.2.2 Posterior Sampling
Table 1: Quantitative results on $1024$ examples of size $256\times 256$ from FFHQ test set. All methods were evaluated under the same experimental setup using LDM. Lower ($\downarrow$) is better in all metrics.
Inpainting (Box) Inpainting (Free-Form) Gaussian Deblurring Super-Resolution (8×8$\times$8 ×)
FID NIQE LPIPS FID NIQE LPIPS FID NIQE LPIPS FID NIQE LPIPS
LDPS 39.81 7.592 0.236 40.17 7.609 0.212 29.30 6.538 0.237 29.64 6.412 0.282
LTDS 39.57 7.602 0.236 39.78 7.578 0.212 30.23 6.553 0.238 30.45 6.412 0.284
ReSample 86.79 7.142 0.230 37.01 6.622 0.151 39.80 7.441 0.275 59.23 7.307 0.356
PSLD 39.68 6.544 0.246 36.26 6.835 0.216 36.31 6.802 0.341 40.33 6.803 0.347
PFLD 39.06 6.509 0.245 36.43 6.817 0.215 37.16 6.751 0.343 38.11 6.832 0.345
LatentDAPS 60.24 9.999 0.257 54.40 8.766 0.223 54.28 9.496 0.283 70.24 10.17 0.344
LD-SMC (1 particle) 33.37 7.032 0.212 33.67 7.034 0.194 29.19 6.575 0.232 30.02 6.426 0.277
LD-SMC (5 particles) 33.87 7.066 0.211 33.60 7.021 0.194 29.47 6.528 0.233 30.62 6.455 0.278
Sequence weights and proposal distributions. To derive an SMC procedure using the proposed target distributions, it is common to use importance sampling. The key idea is to construct proposal distributions $\pi_{t}({\mathbf{z}}_{t:T})$, one for each timestep, from which it is easy to sample, and approximate $p({\mathbf{z}}_{t:T}|{\mathbf{y}}_{0:T})$ with the empirical distribution $\sum_{i=1}^{N}w_{t}^{(i)}\delta_{{\mathbf{z}}_{t:T}^{(i)}}({\mathbf{z}}_{t:T})$. Where, $\delta_{{\mathbf{z}}_{t:T}}$ is the Dirac measure and the weight $w_{t}$ (presented here for $t<T$) is defined as,
$w_{t}$ $\propto\frac{\tilde{p}_{\theta}({\mathbf{z}}_{t:T}|{\mathbf{y}}_{t:T},{\mathbf{y}}_{0})}{\pi_{t}({\mathbf{z}}_{t:T})}$ (5)
$=\frac{\tilde{p}_{\theta}({\mathbf{z}}_{t}|{\mathbf{z}}_{t+1},{\mathbf{y}}_{t},{\mathbf{y}}_{0})}{\pi_{t}({\mathbf{z}}_{t}|{\mathbf{z}}_{t+1})}\frac{\tilde{p}_{\theta}({\mathbf{z}}_{t+1:T}|{\mathbf{y}}_{t+1:T},{\mathbf{y}}_{0})}{\pi_{t+1}({\mathbf{z}}_{t+1:T})}$
$=\frac{\tilde{p}_{\theta}({\mathbf{z}}_{t}|{\mathbf{z}}_{t+1},{\mathbf{y}}_{t},{\mathbf{y}}_{0})}{\pi_{t}({\mathbf{z}}_{t}|{\mathbf{z}}_{t+1})}w_{t+1}.$
Here, $\tilde{p}_{\theta}(\cdot)$ is the unnormalized approximate target density. Plugging the unnormalized target densities from Eq. 4 in Eq. 5 for all timesteps $t$ results in,
$w_{t}\propto\begin{cases}\frac{p_{\theta}({\mathbf{y}}_{T}|{\mathbf{z}}_{T})\bar{p}_{\theta}({\mathbf{y}}_{0}|{\mathbf{z}}_{T})p_{\theta}({\mathbf{z}}_{T})}{\pi_{T}({\mathbf{z}}_{T})},&t=T,\\ \frac{p_{\theta}({\mathbf{y}}_{t}|{\mathbf{z}}_{t})\bar{p}_{\theta}({\mathbf{y}}_{0}|{\mathbf{z}}_{t})p_{\theta}({\mathbf{z}}_{t}|{\mathbf{z}}_{t+1})}{\bar{p}_{\theta}({\mathbf{y}}_{0}|{\mathbf{z}}_{t+1})\pi_{t}({\mathbf{z}}_{t}|{\mathbf{z}}_{t+1})}w_{t+1},&0<t<T,\\ \frac{p_{\theta}({\mathbf{y}}_{0}|{\mathbf{z}}_{0})p_{\theta}({\mathbf{z}}_{0}|{\mathbf{z}}_{1})}{\bar{p}_{\theta}({\mathbf{y}}_{0}|{\mathbf{z}}_{1})\pi_{0}({\mathbf{z}}_{0}|{\mathbf{z}}_{1})}w_{1},&t=0.\end{cases}$ (6)
@@
With these weights, we now turn to defining the proposal distributions $\pi_{t}$ for all timesteps $t$. The optimal choice in the sense of minimizing the variance of the weights is $\pi({\mathbf{z}}_{t}|{\mathbf{z}}_{t:T})=p_{\theta}({\mathbf{z}}_{t}|{\mathbf{z}}_{t+1},{\mathbf{y}}_{0:t})$ (Doucet et al., 2000; Särkkä, 2013). However, they cannot be obtained in closed form. Hence, we design alternative proposal distributions. The proposal distributions for timesteps $t<T$ are defined to be Gaussian $\pi_{t}({\mathbf{z}}_{t}|{\mathbf{z}}_{t+1})=\mathcal{N}({\mathbf{m}}_{t},{\mathbf{S}}_{t})$ with parameters:
${\mathbf{S}}_{t}$ $=\tilde{\sigma}_{t+1}^{2}{\mathbf{I}}$ (7)
${\mathbf{m}}_{t}$ $=\bm{\mathbf{\mu}}_{\theta}({\mathbf{z}}_{t+1},{t+1})-(\gamma_{t}\nabla_{{\mathbf{z}}_{t+1}}\log~{}\bar{p}_{\theta}({\mathbf{y}}_{0}|{\mathbf{z}}_{t+1})$
$~{}~{}~{}+\lambda_{t}\nabla_{\bm{\mathbf{\mu}}_{\theta}({\mathbf{z}}_{t+1},{t+1})}\log~{}q_{\theta}({\mathbf{y}}_{t}|{\mathbf{z}}_{t+1})).$
Where $q_{\theta}({\mathbf{y}}_{t}|{\mathbf{z}}_{t+1})=\mathcal{N}({\mathbf{y}}_{t}|{\mathcal{A}}({\mathcal{D}}(\bm{\mathbf{\mu}}_{\theta}({\mathbf{z}}_{t+1},{t+1}))),\tau^{2}{\mathbf{I}})$, and $\gamma_{t}$, $\lambda_{t}$ are finite scaling coefficients. In addition, $\pi_{T}({\mathbf{z}}_{T})$, the distribution of the proposal at time $T$, is set to the diffusion prior distribution, $\pi_{T}({\mathbf{z}}_{T})=\mathcal{N}(0,{\mathbf{I}})$.
@@
In Eq. 7, we set the variance to be the variance of the prior diffusion model, namely $\tilde{\sigma}_{t+1}=\sigma_{t+1}$; however, other choices are also applicable. The idea behind the proposal mean is to correct the prior mean estimation by shifting it toward latents that agree more strongly with both ${\mathbf{y}}_{t}$ and ${\mathbf{y}}_{0}$. Specifically, the second term of the correction can be seen as making one gradient update step starting from the current prior mean, which serves as an estimate to ${\mathbf{z}}_{t}$.
@@
The parameters $\gamma_{t}$ and $\lambda_{t}$ control the effect of the correction terms to the prior mean. In practice, it is challenging to control the trade-off between the two correction terms. Hence, we propose the following simple approach. During the sampling process, the first term only is used, that is, $\lambda_{t}=0$, and then starting from some predefined timestep $s$, the second term is used as well. The intuition here is that during the initial sampling steps, the quality of the labels ${\mathbf{y}}_{t}$ may not be good. Therefore, we rely on the first term through the posterior mean estimator to capture the general semantics of the image. However, in later sampling stages the quality of the labels increases (see Figure 7 in the appendix), and the latter correction term can help capture fine details in the image. This design choice also relates to the three-stage phenomenon in the diffusion sampling process witnessed in the literature (Yu et al., 2023). Setting $\lambda_{t}=0$ for all timesteps reduces the LD-SMC proposal update to that of TDS (Wu et al., 2024). Appendix D shows an instantiation of $\gamma_{t}$ and $\lambda_{t}$ used in this study.
Having defined the weights and the proposal distributions, we present LD-SMC concisely in Algorithm 1. The algorithm has two parts; the first part is the Gibbs sampling process and the second is the SMC sampling using SISR. In the algorithm the abbreviation ”Multi” refers to Multinomial distribution and a resampling step is performed at each iteration, although it is not mandatory. Appendix C provides a proof that LD-SMC can achieve an arbitrarily accurate estimate of $p_{\theta}({\mathbf{z}}_{0}|{\mathbf{y}}_{0})$ under several conditions. Importantly, the estimate does not depend on any of the approximations made to derive our model. The following Theorem summarizes that informally,
Theorem 4.1.
(informal) Let ${\mathbb{P}}_{N}({\mathbf{z}}_{0:T})=\sum_{i=1}^{N}w^{(i)}_{0}\delta_{{\mathbf{z}}_{0:T}^{(i)}}({\mathbf{z}}_{0:T})$ be the discrete measure obtained by the function $\mathbf{SISR}$ in Algorithm 1. Under regularity conditions ${\mathbb{P}}_{N}({\mathbf{z}}_{0:T})$ converges setwise to $p_{\theta}({\mathbf{z}}_{0:T}|{\mathbf{y}}_{0:T})$ as $N\rightarrow\infty$. Furthermore, the stationary distribution of the Gibbs sampling process is $p_{\theta}({\mathbf{z}}_{0:T},{\mathbf{y}}_{1:T}|{\mathbf{y}}_{0})$, and the marginal $p_{\theta}({\mathbf{z}}_{0}|{\mathbf{y}}_{0})$ is the limiting distribution of the ${\mathbf{z}}_{0}$’s subchain.
Connection to TDS. While the derivation is different, the SMC procedure in LD-SMC resembles somewhat that of TDS (Wu et al., 2024). The main difference between the two methods is the dependence on the auxiliary variables ${\mathbf{y}}_{1:T}$, which are only relevant for LD-SMC in both the proposal distributions and the weights. Empirically, we observed that this additional conditioning helped to better align the sampling with the corrupted image ${\mathbf{y}}_{0}$ compared to using the posterior mean approximation as in (Chung et al., 2023b) and (Wu et al., 2024).
@@ 5 Experiments
Table 2: Quantitative results on $1024$ examples of size $256\times 256$ from ImageNet test set. All methods were evaluated under the same experimental setup using LDM. Lower ($\downarrow$) is better in all metrics.
Inpainting (Box) Inpainting (Free-Form) Gaussian Deblurring Super-Resolution (8×8$\times$8 ×)
FID NIQE LPIPS FID NIQE LPIPS FID NIQE LPIPS FID NIQE LPIPS
LDPS 65.04 7.935 0.379 53.47 7.867 0.334 52.48 6.855 0.383 61.02 6.514 0.439
LTDS 64.74 7.907 0.378 52.75 7.884 0.334 50.82 6.695 0.379 59.12 6.270 0.435
ReSample 90.32 8.464 0.318 44.15 7.104 0.248 46.45 7.411 0.353 87.65 8.290 0.491
PSLD 71.15 8.042 0.434 62.38 8.037 0.411 60.68 6.599 0.417 66.56 7.669 0.489
PFLD 72.83 7.933 0.436 63.34 8.026 0.414 60.94 6.733 0.421 64.72 7.685 0.492
LatentDAPS 98.24 11.36 0.394 68.65 9.625 0.330 77.09 10.55 0.417 104.6 12.68 0.489
LD-SMC (1 particle) 51.49 6.878 0.326 38.21 6.969 0.285 52.48 6.855 0.383 58.06 6.243 0.434
LD-SMC (5 particles) 50.67 6.891 0.325 36.18 6.671 0.278 52.29 6.791 0.383 57.89 6.238 0.433
@@ 5.1 Experimental Setting
We evaluated LD-SMC on ImageNet (Russakovsky et al., 2015) and FFHQ (Karras et al., 2019); both are common in the literature of inverse problems (Chung et al., 2023b; Dou & Song, 2024). In ImageNet, samples were conditioned on the class label. The guidance scale was fixed to $1.0$ in all our experiments. Results can be improved by adjusting it (Rombach et al., 2022). Images were resized to $3\times 256\times 256$ and normalized to the range $[0,1]$. We used the latent diffusion models VQ-4 / CIN256-V2 (Rombach et al., 2022) as the prior model in FFHQ / ImageNet, respectively, with the DDIM diffusion sampler (Song et al., 2021a), according to the data split in (Esser et al., 2021). We sampled $1024$ random images from the validation set of each dataset which were used to evaluate all methods. We followed the protocol of (Song et al., 2024) and added Gaussian noise with zero mean and standard deviation $\tau=0.01$ to the corrupted images. Full experimental details are provided in Appendix A.
@@
Compared methods. We compare LD-SMC with several recent SoTA baseline methods. All methods were evaluated in a similar experimental setup using the same latent diffusion model to ensure fairness in the comparisons. (1) DPS (Chung et al., 2023b), which introduces correction to the sampling process of the diffusion through the posterior mean estimator; (2) TDS (Wu et al., 2024), which uses the twisting technique for approximate sequential Monte Carlo sampling; since DPS and TDS were designed for pixel-space diffusion models, we term the adapted versions of them to the latent space LDPS and LTDS, respectively. (3) ReSample (Song et al., 2024), which applies an optimization procedure during the sampling process to match the approximate posterior mean to the label followed by a resampling step; (4) PSLD (Rout et al., 2023), which introduces a correction term to the DPS step to “glue” ${\mathbf{z}}_{0}$; (5) PFLD (Nazemi et al., 2024), a particle filtering approach for LDM based on PSLD update for the proposal distribution; (6) LatentDAPS which decouples the latents in the sampling trajectory. In all experiments, we evaluated LD-SMC with $N=\{1,5\}$ particles. Using one particle incurs a computational cost similar to that of LDPS (see Table 4 in the appendix). LTDS and PFLD were evaluated with $N=5$ particles. All methods were tested on the following inverse problem tasks: box inpainting, Gaussian deblurring, super-resolution (8×8$\times$8 ×) and free-form inpainting. We followed the setup in (Chung et al., 2023b) for the first three tasks and the setup in (Saharia et al., 2022) for the last one with a few changes to the default hyperparameters. In inpainting tasks, the masks were randomly sampled for each new sample.
@@
Evaluation metrics. In the main text we report the following metrics, FID (Heusel et al., 2017), NIQE (Mittal et al., 2012), and LPIPS (Zhang et al., 2018). Full results including PSNR and SSIM (Wang et al., 2004) are deferred to Appendix H. FID and NIQE are considered perceptual metrics; lower values in them indicate higher perceptual quality. The other metrics can be considered as distortion metrics, which quantify some discrepancy between the generated images and the ground-truth values. Since perceptual metrics and distortion metrics can be in conflict with each other (Blau & Michaeli, 2018), we put more emphasis on perceptual quality. Hence, for all methods, we performed grid search over hyper-parameters and chose the best configuration according to the FID.
(a) Effect of $s$
(b) Effect of $N$
Figure 5: FID and PSNR values when varying $s$ (left) and the number of particles $N$ (right) on ImageNet box (left) and free-form (right) inpainting tasks.
@@ 5.2 Experimental Results
Quantitative results are shown in Tables 1 and 2. From the tables, LD-SMC is usually the best or second best among all the comparisons. Specifically, on inpainting where extrapolation is needed within the masked region and details should be preserved outside of it, LD-SMC can greatly improve over baseline methods. This property is also manifested in Figure 3 and Figure 8 in the appendix. LD-SMC manages to produce plausible reconstructions while maintaining fine details. The differences are especially highlighted on the ImageNet dataset which has more diversity in it. Also, as is clear from the figures, ReSample images suffer from significant artifacts. We speculate that it partly stems from the complete optimization process performed in every few sampling steps according to this method.
Figure 4 shows qualitative results for Gaussian deblurring of images by all methods. From the figure, LD-SMC, LDPS, LTDS, and ReSample are able to generate plausible reconstructions on this task. PSLD and PFLD images have noticeable artifacts, while LatantDAPS images are blurry.
Additional quantitative and qualitative results can be found in Appendix H and I. In terms of the distortion metrics PSNR and SSIM, LD-SMC usually substantially outperforms PSLD and PFLD and has a small advantage over LDPS and LTDS, but overall the three methods are comparable. ReSample and LatentDAPS are usually the best. Nevertheless, as can be seen visually and as indicated by other metrics, it comes at the expense of perceptual quality. LatentDAPS images are overly smooth and details are not preserved in them, while ReSample images have noticeable artifacts, especially in inpainting tasks.